Order: Heparin Sulfate loading dose (bolus) 10,000 units IVP. Following the bolus, administer Heparin Sulfate continuous infusion 15 units/kg/hr.  

Available: Heparin 5,000 u/mL for the loading dose; Heparin 20,000 u/500 mL for the continuous infusion.

Patient weighs 110 pounds.

How many mL's will the nurse administer for the loading dose?

Write only the number in the space provided.

Please show me the work for this problem
Answer #1

ANSWER.

Calculation of amount (ml) administered as loading dose.

        Formula = Desired dose /Dose in hand * Quantity.

          Here

                 Desired dose (D) = 10000 units.

                 Dose in hand (H) = 5000 units.

                       Quantity (Q) = 1ml.

           As per formula

              Amount (Ml) required as loading dose = D/H *Q

                                                                         = 10000 / 5000 * 1

                                                                         = 2 * 1

                                                                         = 2 ml.